OpenStudy (anonymous):

Line x=3 is the axis of symmetry for the graph of a parabola. If the parabola contains the points (1,0) and (4,-3), what is the equation of the parabola? I asked this before, but i got confused again.

OpenStudy (sirm3d):

the equation of the parabola is \[(y-k)=a(x-h)^2\] the line of symmetry is x=3, that makes \[h=3\] use x=1 and y=0 to to obtain one equation in k and a, use x=4 and y=3 to obtain another equation in k and a. solve the system of equations in k and a.

OpenStudy (anonymous):

what are the variables? what is a, k, and h?